Solve the inequalities. Give your answer in interval notation, and indicate the answer geometrically on the real-number line.

a) t + 6 ≤ 2 + 3t

b) 3(2 – 3x) > 4(1 – 4x) 


1
Expert's answer
2022-06-03T05:49:58-0400

a) "t + 6 \u2264 2 + 3t,"

"t-3t\leq2-6," (we have got all the terms with the variable t on the left and the numbers on the right)

"-2t\leq-4," (we have simplified the both sides)

"t\geq\frac{-4}{-2}," (We have divided by the negative number, so we had to change the direction of the inequality)

"t\geq2,"

in interval notation: "t\in[2, \infty),"

the interpretation on the real-number line:




b) "3(2 \u2013 3x) > 4(1 \u2013 4x),"

"6-9x>4-16x,"  (we have removed parentheses)

"-9x+16x>4-6," (we have got all the terms with the variable x on the left and the numbers on the right)

"7x>-2," (we have simplified the both sides)

"x>-\frac{2}{7}," (We have divided by the positive number, so we hadn't to change the direction of the inequality)

in interval notation: "x\in(-\frac{2}{7}, \infty),"

the interpretation on the real-number line: